Status Report

NASA Space Station On-Orbit Status 11 Jul 2003

By SpaceRef Editor
July 11, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 11 Jul 2003

All ISS systems continue to function nominally, except as noted previously or below.  

CDR Yuri Malenchenko performed maintenance on the common power switching unit (BSK) of the Russian segment’s (RS) onboard equipment control system (SUBA), replacing five fuses in a circuit breaker plug with new fuse links, then storing the old fuses as spares.  The activity was supported by tag-up with MCC-M.  [The SUBA controls, monitors, and diagnoses Service Module (SM) systems status.  It operates using sensor output signals and command radio link SM functional outputs, onboard computer system (BVS) units, SM control panels, and system relay outputs.  Its software resides in the SM central computer (TsVM) and terminal computer (TVM).  The BSKs are used to switch electrical power and protect electrical circuits with fuses against overloads.]

Later, Malenchenko reconfigured electronic cables in the SM to connect the Regul-OS antenna feeder unit (AFU) to the antenna feed in a new arrangement.  This task was deferred from 6/24.

FE/SO Ed Lu performed another ppCO2 (carbon dioxide partial pressure) data take from the new CSA-CP (compound specific analyzer-combustion products) brought up on 11P.  (Last time done: 7/3).   [CSA-CP is a toxicology monitoring device for early CO detection/warning and post-fire decontamination.  It also checks for hydrogen chloride (HCl) and other combustion products and contaminants.  Data takes are regularly collected for downlink.]

Malenchenko performed the periodic air sampling in the station, first using the standard Russian AK-1M sampler device in the SM and FGB.  Later, testing for CO (carbon monoxide) levels, he collected SM air samples with the IPD Draeger tubes sampler.

Ed Lu conducted an inventory and inspection of fire ports in the Node, to verify emergency clearance constraints.  [Primary purpose was to video record the fire ports (panel access openings) for ground evaluation of potential blocking in an emergency.  Rearrangement of easily movable items could be done at the discretion of the crew,.]

The Science Officer also took documentary photography for POC (Payload Operations Center) of the MSG (microgravity science glovebox) umbilical connections, for the historical record.  [FOV (field-of-view) data and downlink info was uplinked overnight.]

The CDR concluded the current run of the Molniya-SM/LSO experiment and uninstalled its hardware from SM window #3 for stowage.  [This latest session was started on 7/7.]

After the aborted installation of the electrical heater of the catalyst cartridge (PKF-T) in the SM’s micropurification filtration system (BMP) on 7/2, which apparently did not fit in place, Malenchenko today was scheduled to check the dimensions of the heater unit and tag up with ground specialists via S-band.

Later, Yuri continued the latest maintenance cycle on the RS ventilation system, today removing and replacing the two PS1/PS2 dust collector filters in the FGB air ventilation system, disposing of the old filters (last time done: 6/6).

After a very successful week’s run, the current EarthKAM (Earth Knowledge Acquired by Middle school students) operations came to an end today.  Ed Lu deactivated the system, dismantled the equipment at the Lab nadir window and stowed its components.  [This was the first ever “Summer” mission for EK and the results have reportedly been “terrific”.  In addition to the normal student use, many teachers received hands-on training so they can bring EK into their classrooms.]

Ed supported the ongoing NGL (Next Generation Laptop) testing, which today called for the third data collection and log file take.  [Ed’s periodic support consists of copying  the accumulated files of the automated test program from the new laptop’s microdrive over to the SSC (station  support computer) laptop for data collection and subsequent downlink.  The microdrive is then returned to the NGL and the DOS-based test program restarted.]

Ed Lu completed his ninth weekly filling-out of the FFQ (food frequency questionnaire), which keeps a log of his nutritional intake over time on special medical equipment computer (MEC) software.

Yuri Malenchenko attended to his regular daily inspection of the BIO-5 Rasteniya-2 (“Plants-2”) experiment which studies growth and development of plants (peas) under spaceflight conditions in the Lada-2 greenhouse.

The CDR also completed the regular periodic (weekly) inspection of the active BRPK-1 air/liquid condensate separator in the SM.

The SO performed the periodic transfer of physical exercise data files of the TVIS treadmill, RED resistive exerciser and CEVIS cycle ergometer to the MEC (medical equipment computer), for subsequent downlink to the ground (last time done: 7/3).  [The file transfers involve PCMCIA (portable computer memory card international adapter) cards from TVIS and CEVIS, plus the entries from the RED exercise log book, individually for each crewmember.]

Later,  Ed Lu also copied the accumulated data files from the receiving stations of the wrist-worn HRM (heart rate monitor) to the MEC for downlink, then deleted them on the HRM (last time done: 7/3). 

On the operational PCS (portable computer system) laptops in the ISS, Ed performed the regular once-a-week maintenance reboot.

Yuri prepared the daily inventory management system (IMS) “delta” file for updating the database and conducted the weekly IMS tagup with the ground via S-band, in response to a number of uplinked questions on stowage locations.

Ed Lu took care of the daily status checkup of autonomous Lab payloads (SAMS, MAMS, PCG-STES010).

Also on Ed’s schedule was the third session for his recording of video scenes of a two-day onboard tour for the “Live on ISS” PAO TV program, started on Thursday.  Today, he edited the recordings, which will be downlinked on 7/17 (Thursday next week).  The downlink, with live crew narration, will be aired on NASA TV.

At 10:50am EDT, Yuri Malenchenko participated in a teleconference with school children at TsUP/Moscow in a PAO TV downlink initiated by the RS automated daily time sequencer (SPP).  [The live-TV exchange featured student winners of a Space Science Projects competition from different towns of Russia, assembled at TsUP for the ISS event.]

Later in the day (3:55pm EDT), Ed Lu is scheduled to tag up with science students at Los Alamos University, NM, via S-band/audio.

The crew will also have their weekly teleconference with the ISS Flight Director at MCC-H.

At 2:41pm EDT, the station will maneuver to LVLH +YVV/”barbecue” attitude (vertical y-axis in velocity vector, i.e., facing “top” forward) for six orbits, to allow gathering temperature and power data for the scheduled mid-Beta angle range analysis test.  Afterwards, at 1:36am tomorrow morning, flight attitude will return to XPOP.  [Projected power availability in YVV (autotrack): 24.1 kW from 2B & 4B combined.  XPOP (autotrack w/bias): 22.4 kW from both.]

The crew was asked to keep an eye out for tropical storm (hurricane) Claudette and to record photos or video clips, if possible, for forwarding to PAO.  [MCC-H is watching Claudette as “she” is moving in from the Caribbean towards the Gulf of Mexico, possibly gaining hurricane strength and perhaps later requiring shifting ISS Flight Control from Houston to Moscow.  A decision regarding such control transfer, if required, will in all likelihood not have to be made before Monday morning (7/14).]

In both LVLH and XPOP attitudes, the crew has some good opportunities to view the full Moon over the weekend.  A list with viewing times and angles from the Soyuz DM porthole and DC-1 forward window was uplinked.

Today’s CEO targets, no longer limited in the current LVLH attitude and including the targets of the Lewis & Clark 200-year memorial locations, were Rome, Italy (nadir pass), Lisbon, Portugal (nadir pass), Recife, Brazil (nadir pass), Pomeroy, Washington (LEWIS & CLARK SITE: Nadir pass–looking on the south side of the Snake River [within the second bend of the river upstream of its confluence with the Columbia River]), Lemhi Pass, Mont/Idaho (LEWIS & CLARK SITE:  Nadir pass–shooting mountain range crests [dark forest green, as opposed to lighter, browner river valleys] and the crew should have captured this remote pass over the continental divide), Tropical Storm Claudette (Dynamic event.  Claudette was right of track), Buenos Aires, Argentina (nadir pass), Jarvis Island, Central Pacific (Nadir pass.  Detail of this equatorial atoll’s coral reefs was requested, to complete a global mapping study of these highly productive points in the world ocean), Cairo, Egypt (NIGHT PASS. Nadir view), Tel Aviv, Israel (NIGHT PASS. Nadir view), Amman, Jordan (NIGHT PASS. Nadir view), and Damascus, Syria (NIGHT PASS. Nadir view).
CEO images can be viewed at the websites

See also the website “Space Station Challenge” at

Besides the daily CEO (crew earth observation) target list, the station residents were also provided with times of ISS equatorial night crossings for observing the current equatorial-tropical storm zone.

U.S. and Russian Segment Status (as of 1:50pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (16 amp mode).  Vozdukh CO2 scrubber is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is operating.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode (unit is On).  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 28.1; ppO2 (mmHg) — data invalid; ppCO2 (mmHg) — data invalid.
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 20.0.
  • FGB Cabin:  Pressure (mmHg) — 752; temperature (deg C) — n/a.
  • Node:  Pressure (mmHg) — 740.07; temperature (deg C) — 22.2 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 742.29; temperature (deg C) — n/a; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 742.39; temperature (deg C) — n/a; shell heater temp (deg C) — 23.9, ppO2 (mmHg) — 168.3; ppCO2 (mmHg) — 4.3.
  • PMA-1:  Shell heater temp (deg C) — 23.9
  • PMA-2:  Shell heater temp (deg C) — 21.3.
  • (n/a = data not available)

Propulsion System (PS):

  • Total propellant load available: 3733 kg (8230 lb) as of 7/10  [SM(774) + FGB(2777) + Progress M(182) +Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (Blind/triple-angle mode, non solar-tracking, drag reduction)..
  • SM batteries:  Battery #4 is disconnected (failed 6/16); all other batteries (7) are in “Partial Charge” mode.  Batteries #1, #2 and #3 showing degraded capacity. 
  • FGB batteries:  Battery #3 is offline; battery #2 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-1 is operating; INT-2 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-2 MDM is Off; PL-1 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH -YVV (local vertical/local horizontal = “earth-fixed’: z-axis in local vertical, -y-axis in velocity vector [yaw: -10 deg, pitch: -9.3 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management.
  • Solar Beta angle:  38.5 deg (magnitude increasing).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally [IAC-1 (internal audio controller #1) being analyzed after self-test error.  IAC-2 is prime.]
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #1 with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:16am EDT [= epoch]):

  • Mean altitude — 386.7 km
  • Apogee — 392.3 km
  • Perigee — 381.0 km
  • Period — 92.29 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0008337
  • Orbits per 24-hr. day — 15.60
  • Solar Beta Angle — 38.5 deg (magnitude increasing)
  • Mean altitude loss in last 24 hours — 100 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 26479
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

Did you know that…

…the S-band system transmits 8730 words (2 bytes per word) of data every second?  That’s 1508 Mb (megabytes) of data per day. That means that since S-band has been activated, approximately 1.425 Tb (terabytes) of data have been transmitted to the ground. That equals about 15% of the printed collection of the Library of Congress!

SpaceRef staff editor.